OpenStudy (anonymous):

(2x2 – 23x + 63) ÷ (x – 7) steps please, last one

OpenStudy (anonymous):

bet this one factor too right?

OpenStudy (anonymous):

and you can also guess that one of the factors is \(x-7\) so start with \[2x^2-23x+63=(x-7)(\text{something})\] and see if we can figure out the "something"

OpenStudy (anonymous):

first term of "something" must be \(2x\) because \(2x\times x=2x^2\) and last term must be \(-9\) because \(-9\times -7=63\) so you know it is \[2x^2-23x+63=(x-7)(2x-9)\]

OpenStudy (anonymous):

then as before you get \[\frac{2x^2-23x+63}{x-7}=\frac{(x-7)(2x-9)}{x-7}\] cancel to finish